OpenStudy (anonymous):

is the cubed root of 37 irrational?

OpenStudy (anonymous):

sis needs help

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@hartnn

OpenStudy (paki):

what you guess about it...?

OpenStudy (anonymous):

she thinks that it is irrational.

OpenStudy (paki):

ok, tell her, is she know rational and irrational....?

OpenStudy (anonymous):

she says that a rational number can be turned into a fraction and sometimes irrational numbers can be formed into a fraction too

OpenStudy (anonymous):

@paki

OpenStudy (paki):

sorry for this bad lag here... A rational number is a number that can be expressed as a fraction or ratio. An irrational number cannot be expressed as a fraction.

OpenStudy (paki):

so now what she guess about the answer...?

OpenStudy (paki):

i mean is the cube root of 37 is a rational or irrational...?

OpenStudy (anonymous):

she still thinks that it is irrational lol

OpenStudy (paki):

yeah rite... and please have a look here too... http://math.stackexchange.com/questions/102348/how-can-one-prove-that-the-cube-root-of-9-is-irrational

zepdrix (zepdrix):

Hey gamer :) A value is `rational` if it can be written as a ratio of INTEGERS. Example:\[\Large\rm \sqrt{4}=2\qquad=\quad \frac{2}{1}\]square root of 4 is rational because it can be written as the ratio of 2 and 1. Here is a little tip, any number that is NOT a perfect square will be irrational when you take the square root of it. Same goes for cube roots, is 37 a perfect cube?

zepdrix (zepdrix):

27 is a perfect cube, \(\Large\rm 27=3^3\)

OpenStudy (anonymous):

now she said that if the cubed root of 24.625 is irrational.

OpenStudy (paki):

@zepdrix yeah agree... and what about 37...

OpenStudy (anonymous):

she changed her mind about being irrational

zepdrix (zepdrix):

\[\Large\rm 3^3=27\]\[\Large\rm 4^3=64\] 37 falls somewhere in the middle of those values. It is not a perfect cube, therefore the cube root of 37 is irrational. Changed her mind about what? You are so confusing gamer :( I can't understand what you're saying...

OpenStudy (paki):

actually she want now about 24.625 figure...

OpenStudy (anonymous):

sorry i am her younger bro so i dont know this stuff.

OpenStudy (paki):

ok... open this link for her... each and everything is clearly explained there... http://www.regentsprep.org/regents/math/algebra/AOP1/Lrat.htm

OpenStudy (anonymous):

she just went into an irrational number calculator and found out that the cubed root of 24.625 is irrational.

OpenStudy (paki):

yeah so?

OpenStudy (anonymous):

she said thanks for the xtra information lol

OpenStudy (paki):

pleasure...